OpenStudy (jenniferjuice):

Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $21.95 per month plus an additional $0.07 per minute. Cell Best has a plan that costs $25.95 per month plus an additional $0.05 per minute. How many minutes can Sasha talk and have the same cost each month?

OpenStudy (jenniferjuice):

0.12m = 4.00 21.95 + 0.07m = 26.00 21.95 + 0.07m = 25.95 + 0.05m 25.95 + 0.05m = 22.02

OpenStudy (anonymous):

21.95 + .07m = 25.95 + .05m